How many photons are in the laser pulse?

A laser pulse with wavelength 515 nm contains 4.85 mJ of energy.

How many photons are in the laser pulse?
Answer #1

1.24*10^16

source: me
answered by: dd
Answer #2

E = n*h*ν (frequency) , n = E/(h*v)

find the frequency which equals 5.83*10^14s (c/λ)

then h * v = 6.26*10^-34 J/s * 5.83*10^14s = 3.86*10^-19J

lastly E/(h*v) = 1.87*10^16J